OpenStudy (anonymous):

write y-3=-1 (x+7) in standard form

OpenStudy (anonymous):

Multiply −1 by each term inside the parentheses. y−3=−x−7 Since −3 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3 to both sides. y=3−x−7 Simplify the right-hand side of the equation. y=−x−4